OpenStudy (anonymous):

find absolute max or absolute min, f(x,y)=6-x^2-4y^2 with constrain (x,y):-2

OpenStudy (anonymous):

\[f(x,y)=6-x^2-4y^2 ; R{(x,y):-2 \le x \le2, -1 \le y \le 1}\]

zepdrix (zepdrix):

 It always helps to sketch the boundaries of our region :D